OpenStudy (anonymous):

Solve the equation. |-15| + 82 = 80y - |-37| - 79y

OpenStudy (gorv):

mode makes negative value positive 15+82=80y-37-79y=y-37 97+37=y y=134

OpenStudy (gorv):

@MyCatWasHere

OpenStudy (anonymous):

When it has the lines around a number it means the opposite like |-12| is 12 but |12| is 12

OpenStudy (gorv):

it is called mode and it makes the value positive always

OpenStudy (gorv):

@HANNAH89

OpenStudy (anonymous):

Oh okay thanks